# Geometry-Induced Localization and Multifractality in Spiral quasiperiodic chain

### v1

@@ I Introduction
Quasiperiodic systems play a key role in modern condensed matter physics, optics, and engineered wave systems. They host a range of interesting phenomena, such as fractal energy spectra, self-dual criticality, mobility edges, localization transitions, and topological boundary states [2, 3, 4, 5]. Among these, the Aubry-André (AA) model is a notable one-dimensional quasiperiodic lattice. It exhibits an exact localization transition at a specific quasiperiodic strength, along with a multifractal spectrum at the self-dual point [2, 4]. This straightforward nature, combined with its rich mathematical framework, has made the AA model essential for studying localization and criticality in both theoretical and experimental settings.
In the last ten years, improvements in experimental techniques have allowed researchers to directly create AA-type quasiperiodic lattices in various physical platforms. Ultracold atoms in bichromatic optical lattices have shown Anderson localization, AA criticality, and mobility-edge behavior, all with single-site resolution [7, 8, 9]. Photonic waveguide arrays have offered additional realizations with strong geometric tunability, which has led to localization, topological pumping, and quasiperiodic boundary states [12, 13, 14]. Acoustic and mechanical metamaterials have also expanded the study of quasiperiodic physics into classical systems, where phononic and mechanical excitations mimic electronic quasiperiodicity [10, 24]. These advances underline that quasiperiodic models are now not only of theoretical significance but also serve as adaptable platforms for investigating complex wave dynamics across different fields.
At the same time, it also become quite clear that geometry and curvature, have a significant impact on quantum and wave phenomena. Curved or strained lattices can create effective gauge fields, change tunneling amplitudes, shift localization transitions, and produce entirely new spectral structures [29, 30, 31]. Helical and spiral geometries, in particular, have gained much attention because of their relevance to experimentally accessible platforms. In photonics, femtosecond-laser-written waveguides can be designed along arbitrary three-dimensional paths, enabling the creation of helical Floquet lattices, synthetic gauge fields, and curvature-induced band engineering [18, 16, 17]. In ultracold atoms, optical vortex beams and Laguerre-Gaussian modes have been employed to form spiral or ring-trap geometries with complex curvature and torsion [34, 35]. At the nanoscale, chiral molecular structures like DNA and helical polymers show geometry-induced transport behaviors, including chiral-induced spin selectivity [40, 41]. These cases highlight that geometric embedding is an effective and experimentally viable way to manipulate the effective hopping landscape of lattice systems. Motivated by these insights, we investigate how introducing a quasiperiodic AA chain into a spiral geometry changes its spectral, localization, and multifractal properties. In our setup, the onsite quasiperiodic potential maintains its typical AA structure, while the curved embedding leads to a systematic change in intersite distances and, therefore, the hopping amplitudes. This creates a geometry-induced inhomogeneous tight-binding model where deviations from standard AA behavior occur only due to curvature effects. We demonstrate that this nonlinear geometric modulation boosts localization at much lower quasiperiodic strengths, reconfigures the overall energy spectrum, and significantly modifies the multifractal nature of the eigenmodes. These effects are similar to those seen in strained and curved photonic systems, but here they result from a controlled geometric embedding. Our findings position spiral geometries as a straightforward, general, and experimentally accessible way to tailor quasiperiodicity, allowing geometric control over localization and multifractality in photonic waveguides, ultracold atoms, mechanical metamaterials, and nanoscale chiral structures.
@@ II.2 Quasiperiodic modulation
where $q$ is a real number and $\tau_{q}$ is the corresponding mass exponent. Extended states satisfy $\tau_{q}=d(q-1)$ with $d=1$, which aligns with the ballistic spreading observed in ultracold-atom and photonic experiments on quasiperiodic lattices [7, 8]. Localized states, on the other hand, have $\tau_{q}=0$. Multifractal states, known to emerge at the localization transition of the Aubry–André model and seen directly in photonic waveguide measurements [12, 47], show a nonlinear dependence $\tau_{q}=D_{q}(q-1)$, where $D_{q}$ represents the generalized fractal dimension. Here, $D_{q}=1$ indicates an extended state, $D_{q}=0$ indicates a localized state, and values in between show a nonuniform fractal distribution of amplitudes.
@@
Panel (a) shows how $\ln(\mathrm{IPR}_{2})$ scales with $\ln(L)$ for three representative eigenstates. A slope near one corresponds to $D_{2}\approx 1$, indicating an extended state with amplitude spreading uniformly across the entire lattice. Such states have been observed in ultracold-atom and photonic realizations of quasiperiodic systems [7, 12]. A slope between zero and one yields $0<D_{2}<1$, showing a multifractal wavefunction where the amplitude is concentrated on a sparse yet significant set of lattice sites. This behavior marks criticality in quasiperiodic lattices and has been experimentally detected in laser-written photonic quasicrystals [19, 14]. Finally, a nearly flat line indicates $D_{2}\approx 0$, typical of a fully localized state confined to just a few sites.
Panel (b) expands on this characterization by showing the generalized scaling $\ln(\mathrm{IPR}_{q})$ for $q=2$ to $5$. The fact that the extracted mass exponents $\tau_{q}$ do not fall on a straight line as a function of $q$ shows real multifractal behavior rather than a simple shift between extended and localized phases. This nonlinear $\tau_{q}$$q$ dependence is a hallmark of intermediate critical states, consistent with both theoretical predictions and experimental findings in quasiperiodically modulated photonic lattices and synthetic matter platforms [19, 14].

### v2

@@ I Introduction
Quasiperiodic systems constitute a central topic in condensed matter physics and wave dynamics, bridging crystalline order and randomness. Since the seminal works on Anderson localization in disordered media [1] and the formulation of quasiperiodic tight-binding models [3, 2], it has been understood that quasiperiodicity can generate rich spectral and transport phenomena distinct from both periodic and random systems. These include fractal energy spectra, critical self-dual points, mobility edges, localization transitions, and multifractal eigenstates [4, 5, 6].
Among such models, the Aubry–André (AA) chain occupies a paradigmatic role. It exhibits an exact metal–insulator transition driven by the strength of a quasiperiodic onsite potential, together with critical multifractal behavior at the self-dual point [2, 4, 7]. The AA model and its generalizations have therefore become standard platforms for exploring the interplay between localization, topology, and quasiperiodicity [8, 9, 10].
Rapid experimental progress over the last decade has enabled direct realizations of quasiperiodic lattices in a wide variety of physical settings. In ultracold atomic gases, bichromatic optical lattices have demonstrated Anderson localization and AA criticality with high controllability [11, 12]. Related phenomena have also been observed in driven and modulated cold-atom systems, including quantized pumping and topological transport [9, 10]. Photonic waveguide arrays provide a particularly flexible platform, enabling the observation of localization, topological pumping, and quasiperiodic boundary states [13, 15, 14, 16]. Comprehensive overviews of discrete photonic lattices and light transport in such media are given in [17].
Beyond optics and cold atoms, quasiperiodic concepts have been extended to classical wave systems, including acoustic and mechanical metamaterials, where phononic and elastic excitations emulate electronic quasiperiodicity [24, 18].Plasmonic and moiré structures provide yet another arena in which geometry and long-range Coulomb interactions give rise to nontrivial collective excitations, including topological and quasiperiodic plasmon modes [30, 29, 28, 31].
In parallel, the role of geometry and curvature in quantum and wave systems has attracted growing interest. Curved and strained lattices can generate effective gauge fields, modify hopping amplitudes, and reshape spectral and topological properties, as demonstrated in strained graphene and related Dirac materials where lattice deformations act as synthetic gauge fields [33, 34, 35]. More generally, geometric and quantum metric effects have been shown to influence transport and topological responses in crystalline systems [36]. Such geometric effects are experimentally accessible in several platforms, including photonic lattices and other synthetic systems where lattice geometry and curvature can be precisely engineered [16, 32]. In photonics, three-dimensional waveguide architectures enable helical and Floquet lattices with engineered curvature [17]. In ultracold atoms, optical vortex and Laguerre–Gaussian beams allow the realization of ring and spiral trapping geometries [20, 27, 21, 32].
At the nanoscale, chiral and spiral molecular structures introduce geometry-driven transport phenomena, including chiral-induced spin selectivity and spin-dependent localization [23, 26]. These systems highlight how structural chirality can strongly influence electronic and spin dynamics.
Motivated by these developments, we explore how embedding a quasiperiodic Aubry–André chain into a spiral geometry modifies its spectral, localization, and multifractal properties. While the onsite quasiperiodic modulation retains the standard AA form, the curved embedding induces systematic variations in intersite distances and hence in the hopping amplitudes, yielding a geometry-controlled inhomogeneous tight-binding model. We demonstrate that curvature alone can significantly enhance localization at reduced quasiperiodic strengths, reshape the energy spectrum, and alter the multifractal characteristics of the eigenstates. Our results establish spiral geometries as a simple and experimentally viable route to engineer quasiperiodicity and to control localization and multifractality across photonic, atomic, mechanical, and nanoscale chiral platforms.
@@ II.2 Quasiperiodic modulation
where $q$ is a real number and $\tau_{q}$ is the corresponding mass exponent. Extended states satisfy $\tau_{q}=d(q-1)$ with $d=1$, which aligns with the ballistic spreading observed in ultracold-atom and photonic experiments on quasiperiodic lattices [11, 12]. Localized states, on the other hand, have $\tau_{q}=0$. Multifractal states, known to emerge at the localization transition of the Aubry–André model and seen directly in photonic waveguide measurements [13], show a nonlinear dependence $\tau_{q}=D_{q}(q-1)$, where $D_{q}$ represents the generalized fractal dimension. Here, $D_{q}=1$ indicates an extended state, $D_{q}=0$ indicates a localized state, and values in between show a nonuniform fractal distribution of amplitudes.
@@
Panel (a) shows how $\ln(\mathrm{IPR}_{2})$ scales with $\ln(L)$ for three representative eigenstates. A slope near one corresponds to $D_{2}\approx 1$, indicating an extended state with amplitude spreading uniformly across the entire lattice. Such states have been observed in ultracold-atom and photonic realizations of quasiperiodic systems [11, 13]. A slope between zero and one yields $0<D_{2}<1$, showing a multifractal wavefunction where the amplitude is concentrated on a sparse yet significant set of lattice sites. This behavior marks criticality in quasiperiodic lattices and has been experimentally detected in laser-written photonic quasicrystals [amaral2018synthetic, 14]. Finally, a nearly flat line indicates $D_{2}\approx 0$, typical of a fully localized state confined to just a few sites.
Panel (b) expands on this characterization by showing the generalized scaling $\ln(\mathrm{IPR}_{q})$ for $q=2$ to $5$. The fact that the extracted mass exponents $\tau_{q}$ do not fall on a straight line as a function of $q$ shows real multifractal behavior rather than a simple shift between extended and localized phases. This nonlinear $\tau_{q}$$q$ dependence is a hallmark of intermediate critical states, consistent with both theoretical predictions and experimental findings in quasiperiodically modulated photonic lattices and synthetic matter platforms.